OpenStudy (anonymous):

Solve the system of equations: x-y=4 x-2y=10

OpenStudy (anonymous):

from eqn 1, x = 4+y plugging in the value of x to eqn 2 4+y-2y=10 -y=6 y=-6 and x = 4-6 x=-2

hero (hero):

y = x -4 y = (10-x)/-2 y = y -2(x-4) = 10 -x -2x+8 = 10 -x x = -2 y= -6

OpenStudy (anonymous):

thanks

OpenStudy (lilg132):

use substitution rearrange first formula to either get x = or y = lets choose x = so you would get x = 4 + y now use this in your second equation everywhere you see x just replace it with that so your second equation x - 2y = 10 replace the x with 4 + y so 4 + y - 2y = 10 bring all the numbers to one side and variables on the other remember bringing a number to the otherside changes its value e.g. briging 4 to the other side will change it to - 4 so y - 2y = 10 - 4 is y = 6

myininaya (myininaya):

thats one way you could multiply bottom equation by -1 (or top) and then add the equations

OpenStudy (lilg132):

dont forget to click good answer

myininaya (myininaya):

x - y = 4 -x + 2y = -10 ------------------- 0 +y=-6 =>y=-6

myininaya (myininaya):

and we can plug in y=-6 into either equation given and solve for x

OpenStudy (lilg132):

sorry the answer is y = -6 i made a mistake if there was a edit feature i would have changed that but otherwise i have to rewrite whole thing again y - 2y = 10 - 4 will give you -y = 6 flip them around remember the rule about bringing something to the other side changes their value to the opposite so 6 = y

OpenStudy (lilg132):

arghhh made it again the mistake !! -6 = y

myininaya (myininaya):

\[x -(-6)=4 => x+6=4 => x=4-6=-2\]